OpenStudy (anonymous):

Medal or fan f(x) = x2 + 4, find f(6). 16 36 2 40

OpenStudy (anonymous):

(6)^2 + 4 36+4=40

OpenStudy (anonymous):

Plugging in 6 we get f(6) = 6^2 + 4 f(6) = 40